OpenStudy (anonymous):

Verify that f(x) and f^-1(x) are inverse functions f(x)=10x, f^-1(x)=x/10

OpenStudy (anonymous):

f(x)=10x y=10x x=y/10 then change the plce of y and x after that you have your inverse function that is y=x/10 f(x)-1=x/10

OpenStudy (anonymous):

so both of them are inverse?

OpenStudy (anonymous):

no the second function is the inverse of the first function

OpenStudy (anonymous):

so only one of them? are you positive?

OpenStudy (anonymous):

no each of the is the inverse of the other one

OpenStudy (anonymous):

we have \[\left(f\circ f^{-1}\right)(x)=x\\ {(10x)\over10}=x\]

OpenStudy (anonymous):

QED